OpenStudy (anonymous):

Find the common ratio and a formula for the nth term of the geometric sequence: 1/18, 1/3, 2, 12, 72,... a_n= ?

OpenStudy (anonymous):

Common ratio is 6 because 1/3/1/18 = 18/3 = 6 \[a _{n} = 1/18(6)^{n - 1}\]

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

You're welcome :)

OpenStudy (maheshmeghwal9):

http://www.youtube.com/watch?v=pXo0bG4iAyg

OpenStudy (maheshmeghwal9):

good for knowledge:)